a basket contains 110 oranges and 50 lemons only. How many grape fruits must be added to basket so that the probability of randomly picking a lemon is equal to 0.25
Answer by htmentor(1343)

The probability of picking a lemon is:
P(L) = 50/(110 + 50 + x) = 0.25, where x = the number of grapefruits
Solve for x:
50 = 0.25(160 + x)
50 = 40 + 0.25x
x = 40 grapefruits